# First moments of ${\rm GL}(3)\times{\rm GL}(2)$ and ${\rm GL}(2)$ $L$-functions and their applications to subconvexity

### v1

@@ 3.1.2. Oscillatory case
$\begin{split}&L\geq\frac{M^{1+\varepsilon}}{\sqrt{\mathcal{X}}},\quad L\leq\frac{\mathcal{X}^{1/3-\varepsilon}}{q^{5/3}}\\ &L\leq\frac{\mathcal{X}^{1-\varepsilon}}{q^{2}M},\quad L\leq\left(\frac{\mathcal{X}}{q^{2}}\right)^{1/8-\varepsilon};\end{split}$ (3.52)
while, the second line requires that
$\begin{split}&L\leq\frac{\mathcal{X}^{1/7-\varepsilon}M^{2/7}}{q^{12/7}},\quad L\leq\left(\frac{\mathcal{X}}{q^{6}}\right)^{1/4-\varepsilon}\quad L\leq\frac{\mathcal{X}^{1/5-\varepsilon}M^{2/5}}{q^{8/5}}.\end{split}$ (3.53)
Recall that $(M^{3/2}q^{2})^{1-2\delta}\leq\mathcal{X}\leq(M^{3/2}q^{2})^{1+2\varepsilon}$ for any $\delta,\varepsilon>0$. Collecting all the inequalities (3.25), (3.26), (3.48), (3.52) and (3.53), we infer that the effective range for $L$ should be
$\frac{q^{2+\varepsilon}}{{\sqrt{M}}}+\frac{\sqrt{M}}{q^{2-\varepsilon}}+\frac{M^{1/4+\varepsilon}}{q}+1\leq L\leq\min{\left\{M^{3/16-\varepsilon},\,\frac{\sqrt{M}}{q^{10/7+\varepsilon}},\,\frac{M^{3/8-\varepsilon}}{q}\right\}}.$ (3.54)
This yields the range $M^{5/32+\varepsilon}\leq q\leq M^{7/32-\varepsilon}$, and thus gives the constraint as shown in (1.3). The assertion of Corollary 1.3 now follows by combining with (3.2).
4. Proofs of Theorem 1.4 and Corollary 1.5
In this section, we shall be dedicated to the proofs of Theorem 1.4 and Corollary 1.5.
@@
We shall consider the average of the first moments of $\rm\text{GL}(3)\times GL(2)$ and $\text{GL}(2)$ $L$-functions in a family with a ‘harmonic’ average of $k\sim K$. To be precise, let $W$ be a smooth function compactly supported on $[1/2,5/2]$ with bounded derivatives. Let $f$ be a $\text{GL}(2)$-newform of weight $k\in 2\mathbb{Z}$ and full level. Let $\ell\geq 1$ be an integer. Let $K$ be a sufficiently large parameter, and $K^{100\varepsilon}\ll|t|\ll K^{2-100\varepsilon}$. We attempt to obtain an upper-bound for
$\sum_{k\equiv 0\bmod 2}W{\left(\frac{k-1}{K}\right)}\sum_{f\in\mathcal{B}^{\ast}_{k}}\omega^{-1}_{f}\lambda_{f}(\ell)L(1/2+it,F\otimes f).$ (4.55)
As described in §2.3, the root number $\varepsilon(F\otimes f)$ is dependent on the weight $k$; thus, by the function equation method, we need to turn to seeking an effective control over the following two sums
$\sum_{k\equiv 0\bmod 2}W{\left(\frac{k-1}{K}\right)}\sum_{f\in\mathcal{B}^{\ast}_{k}}\omega^{-1}_{f}\lambda_{f}(\ell)\,\sum_{n\geq 1}\frac{A_{F}(n,1)\lambda_{f}(n)}{n^{1/2+it}}V^{\star}{\left(\frac{n}{\mathcal{Y}}\right)},$ (4.56)
@@
$\sum_{k\equiv 0\bmod 2}i^{k}\,W{\left(\frac{k-1}{K}\right)}\sum_{f\in\mathcal{B}^{\ast}_{k}}\omega^{-1}_{f}\lambda_{f}(\ell)\,\sum_{n\geq 1}\frac{A_{F}(n,1)\lambda_{f}(n)}{n^{1/2+it}}V{\left(\frac{n}{\mathcal{Y}}\right)},$ (4.57)
where $(|t|k)^{3-\delta}\leq\mathcal{Y}\leq(|t|k)^{3-\varepsilon}$ for arbitrarily small constants $\delta,\varepsilon>0$, and $V^{\star},V$ are two smooth bump functions supported $[1/2,5/2]$, identical to one for $x\in[1,2]$, with derivatives satisfying ${V^{\star}}^{(j)},V^{(j)}\ll_{j}1$ for any $j\geq 0$. We shall merely handle the second sum (4.57) in the following. The same analysis works for the first sum (4.56). The main difference is that, for the first sum, the Lemma 2.8 will be used, which, however, produces a possibly weaker bound since of more error terms thereof; for the second one, we will apply Lemma 2.9 instead after a while.
By equipping with Petersoon formula, Lemma 2.4, we split the expression in (4.57) into the diagonal term and the off-diagonal term. The diagonal term is roughly $\ll K^{1+\varepsilon}/\sqrt{\ell}$; while, for the off-diagonal term, it is presented as the following form
$2\pi\sum_{c\geq 1}\sum_{k\equiv 0\bmod 2}W{\left(\frac{k-1}{K}\right)}\sum_{n\geq 1}\frac{A_{F}(n,1)}{n^{1/2+it}}\frac{S(n,\ell;c)}{c}J_{k-1}{\left(\frac{4\pi\sqrt{n\ell}}{c}\right)}V{\left(\frac{n}{\mathcal{Y}}\right)}.$
@@
$\mathcal{T}(K,t)=\sum_{c\ll\sqrt{\mathcal{Y}\ell}/K^{2-\varepsilon}}\,\sum_{n\geq 1}\frac{A_{F}(n,1)}{n^{1/2+it}}\frac{S(n,\ell;c)}{c}\,e{\left(\frac{4\pi\sqrt{n\ell}}{c}\right)}V{\left(\frac{n}{\mathcal{Y}}\right)},$ (4.58)
@@
$\widehat{\mathcal{T}}(K,t)=\sum_{\sqrt{\mathcal{Y}\ell}/K^{1-\varepsilon}\ll c\ll\sqrt{\mathcal{Y}\ell}/K^{1-100\varepsilon}}\,\sum_{n\geq 1}\,\frac{A_{F}(n,1)}{n^{1/2+it}}\frac{S(n,\ell;c)}{c}W{\left(\frac{\sqrt{n\ell}}{cK}\right)}V{\left(\frac{n}{\mathcal{Y}}\right)}$ (4.59)
plus two error terms $\text{Err}^{\flat}(K,t)$ and $\text{Err}^{\natural}K,t)$, say, which can be estimated as
$\begin{split}\text{Err}^{\flat}(K,t)&\ll\frac{1}{K^{2-\varepsilon}}\sum_{n\sim\mathcal{Y}}\frac{|A_{F}(n,1)|}{\sqrt{n}}\sum_{c\sim\sqrt{\mathcal{Y}\ell}/K}\frac{|S(n,l;c)|}{c}\ll\frac{X^{3/4}\ell^{1/4}}{K^{5/2-\varepsilon}},\end{split}$ (4.60)
@@
$\begin{split}\text{Err}^{\natural}(K,t)&\ll\frac{\sqrt{\ell}}{K^{4-\varepsilon}}\sum_{n\sim\mathcal{Y}}{|A_{F}(n,1)|}\sum_{c\ll\sqrt{\ell}\mathcal{Y}^{1/2+\varepsilon}}\frac{|S(n,l;c)|}{c^{2}}\ll\frac{\mathcal{Y}\sqrt{\ell}}{K^{4-\varepsilon}},\end{split}$ (4.61)
respectively. Here, $\text{Err}^{\flat}(K,t)$ comes from the first error term in (2.24); while, $\text{Err}^{\natural}(K,t)$ is for the second one. We have applied bound (2.20) in estimating $\text{Err}^{\natural}(K,t)$, when $k$ is sufficiently large.
In the following paragraphs, we will be devoted to the estimates of $\mathcal{T}$ and $\widehat{\mathcal{T}}$ given as in (4.58) and (4.59), respectively.
4.1.1. Bounding $\mathcal{T}$
We first take care of $\mathcal{T}$. By the dyadic subdivision, what we need to be concentrated on actually is the following
$\sup_{C\ll\sqrt{\mathcal{Y}\ell}/K^{2-\varepsilon}}\sum_{c\sim C}\frac{1}{c}\sum_{n\geq 1}\frac{A_{F}(n,1)}{\sqrt{n}}S(n,\ell;c)V_{c,\ell}{\left(\frac{n}{\mathcal{Y}}\right)},$ (4.62)
with
$V_{c,\ell}(x)=V(x)\,e{\left(-\frac{t\log x\mathcal{Y}}{2\pi}+\frac{4\pi\sqrt{x\mathcal{Y}\ell}}{c}\right)}$
for any $x\in\mathbb{R}^{+}$.
Case I: when $C$ is small. Let $C^{\star}\geq 1$ be a parameter determined later. We first consider the case where $C$ is suitably small, such that $C<C^{\star}$. In this case, it is clear that
$\mathcal{T}(K,t)\ll{C^{\star}}^{1/2+\varepsilon}\sqrt{\mathcal{Y}}$ (4.63)
by using the Weil bound for Kloosterman sums.
Case II: when $C$ is large. Now, we turn to the case where $C\geq C^{\star}$. By invoking the Voronoĭ summation formula, Lemma 2.1, this sum is thus essentially turned into
$\begin{split}\sup_{C^{\star}\leq C\ll\sqrt{\mathcal{Y}\ell}/K^{2-\varepsilon}}\frac{C}{\sqrt{\mathcal{Y}}}&\sum_{c\sim C}\sum_{\begin{subarray}{c}\pm,m^{\ast},r\\ m^{\ast}\geq 1\\ r|c\end{subarray}}\frac{\overline{A_{F}(m^{\ast},r)}}{m^{\ast}r^{2}}\,e\left(\mp\frac{m^{\ast}r^{2}\overline{\ell}}{c}\right)\,\Omega_{\pm}\left({\frac{m^{\ast}r^{2}\mathcal{Y}}{c^{3}};{V}_{c,\ell}}\right).\end{split}$ (4.64)
It suffices to consider the salient scenario where $r=1$, as hinted by the preceding argument in §3.2; the disturbance of the parameter $r$ has no any influence on the final result by an adaptation of the same analysis. The heart of the task thence lies on
$\sup_{C^{\star}\leq C\ll\sqrt{\mathcal{Y}\ell}/K^{2-\varepsilon}}\frac{C}{\sqrt{\mathcal{Y}}}\sum_{c\sim C}\sum_{\begin{subarray}{c}\pm,m^{\ast}\geq 1\end{subarray}}\frac{\overline{A_{F}(m^{\ast},r)}}{m^{\ast}}e\left(\mp\frac{m^{\ast}\overline{\ell}}{c}\right)\,\Omega_{\pm}\left({\frac{m^{\ast}\mathcal{Y}}{c^{3}};{V}_{c,\ell}}\right),$
which, by Lemma 2.2, is approximately
$\begin{split}\sup_{C^{\star}\leq C\ll\sqrt{\mathcal{Y}\ell}/K^{2-\varepsilon}}\frac{\mathcal{Y}^{1/6}}{C}&\sum_{c\sim C}\sum_{\begin{subarray}{c}\pm,m^{\ast}\geq 1\end{subarray}}\frac{\overline{A_{F}(m^{\ast},1)}}{{m^{\ast}}^{1/3}}\,e\left(\mp\frac{m^{\ast}\overline{\ell}}{c}\right)\,\widehat{V_{\pm,\ell}}\left({\frac{m^{\ast}\mathcal{Y}}{c^{3}}}\right),\end{split}$ (4.65)
@@ 4.1.1. Bounding $\mathcal{T}$
$\widehat{V_{\pm,\ell}}(x)=\int_{\mathbb{R}^{+}}V(y)\,e{\left(-\frac{t\log y}{2\pi}+\frac{4\pi\sqrt{y\mathcal{Y}\ell}}{c}\pm 3(xy)^{1/3}\right)}\mathrm{d}y.$
Lemma 4.1.
Let the integral $\widehat{V_{\pm,\ell}}(x)$ be defined as above. Then we have
  • $\widehat{V_{+,\ell}}(x)\ll\mathcal{Y}^{-A}$ for any sufficiently large $A>0$;
  • $\widehat{V_{-,\ell}}(m^{\ast}\mathcal{Y}/c^{3})$ is negligibly small unless that $m^{\ast}\asymp\sqrt{\mathcal{Y}}\ell^{3/2}$, and in which case it is essentially given by
    $\widehat{V_{-,\ell}}\left({\frac{m^{\ast}\mathcal{Y}}{c^{3}}}\right)\asymp\frac{\sqrt{c}}{(\mathcal{Y}m^{\ast})^{1/6}}\,e{\left(\frac{2\mathcal{Y}^{1/6}{m^{\ast}}^{2/3}}{c\sqrt{\ell}}+O{\left(\frac{(|t|+1){m^{\ast}}^{1/3}}{\mathcal{Y}^{1/6}\sqrt{\ell}}\right)}\right)}.$ (4.66)
Proof.
For any 1/$21/2\leq y\leq 5/$ 2, one defines
$f(y)=\mathcal{A}y^{1/2}\pm\mathcal{B}y^{1/3}-{t}\log y/{2\pi},$ (4.67)
with $\mathcal{A}:=\sqrt{\mathcal{Y}\ell}/c$, and $\mathcal{B}:=3x^{1/3}$. Notice that $\mathcal{A}\geq K^{2}\gg t$. One thus finds that the $+^{\prime}$ case will become negligible, by repeated integration by parts many times. Next, we shall mermly consider the $-^{\prime}$ case, and compute the root of $f^{\prime}(y)=0$; this yields
$\frac{1}{2}\mathcal{A}y^{1/2}-\frac{1}{3}\mathcal{B}y^{1/3}-t/2\pi=0.$ (4.68)
We may approximate the root $y=y^{\flat}$ by writing $y^{\flat}=y_{0}+y_{1}+y_{2}+\cdots$. To be specific, set $y_{0}=\left({2\mathcal{B}}/{3\mathcal{A}}\right)^{6}$ (which follows by equating the first and second terms on the left-hand side of the equation). One computes $y_{1}$ from the equation
$\frac{1}{2}\mathcal{A}\left(y^{1/2}_{0}+\frac{1}{2}y_{1}y^{-1/2}_{0}\right)-\frac{1}{3}\mathcal{B}y^{1/3}_{0}-t/2\pi=0,$
where we have approximated $y^{1/2}$ by $y^{1/2}_{0}+y_{1}y^{-1/2}_{0}/2$, and $y^{1/3}$ by $y^{1/3}_{0}$ in (4.68), respectively. This gives that $y_{1}=\frac{2t}{\pi\mathcal{A}}\sqrt{y_{0}}$. Further, by solving the equation
$\frac{1}{2}{\mathcal{A}}\left(y^{1/2}_{0}+\frac{1}{2}(y_{1}+y_{2})y^{-1/2}_{0}\right)-\frac{1}{3}\mathcal{B}\left(y^{1/3}_{0}+\frac{1}{3}y_{1}y^{-2/3}_{0}\right)-t/2\pi=0,$
we infer that $y_{2}=\frac{\mathcal{B}}{\mathcal{A}}\frac{8t}{9\pi\mathcal{A}}y^{1/3}_{0}$. It follows that $\widetilde{V}_{-,\ell}(x)$, as a function with respect to $x$, is essentially given by the following
$\frac{1}{{\mathcal{B}}^{1/2}}{y}^{-it}_{0}\,e\left(\frac{1}{2}{\mathcal{B}}(y_{0}+y_{1}+y_{2}+\cdots)^{1/3})\right)\approx\frac{1}{{\mathcal{B}}^{1/2}}{y}^{-it}_{0}\,e\left(\frac{2\mathcal{B}^{2}}{9\mathcal{A}}+O{\left(\frac{t{\mathcal{B}}}{\mathcal{A}}\right)}\right).$
Note that we have not taken into account the contribution $e(-t\log y_{0}/2\pi)$ in (4.67), as far as the disturbance is concerned. ∎
As a direct consequence, by Lemma 4.1, one finds that the sum (4.65) is essentially of the fowling form
$\begin{split}&\sup_{C^{\star}\leq C\ll\sqrt{\mathcal{Y}\ell}/K^{2-\varepsilon}}\sum_{c\sim C}\frac{1}{\sqrt{c}}\sum_{m^{\ast}\geq 1}\frac{\overline{A_{F}(m^{\ast},1)}}{\sqrt{m^{\ast}}}\,e\left(\frac{m^{\ast}\overline{\ell}}{c}\right)\,\widetilde{{V}}_{c,\ell}\left({\frac{m^{\ast}}{\sqrt{\mathcal{Y}}\ell^{3/2}}}\right),\end{split}$
with $\widetilde{{V}}_{c,\ell}(x)=V(x)e{\left(2\sqrt{\mathcal{Y}\ell}/c\cdot x^{2/3}+O_{c,\ell}(|t|x^{1/3})\right)}$ for any $x\in\mathbb{R}^{+}$. Note that the phase of $\widetilde{{V}}_{c,\ell}$ is roughly $\sqrt{\mathcal{Y}\ell}/c$. One also finds that $\frac{\partial^{i}}{\partial x^{i}}\widetilde{{V}}_{c,\ell}(x)\ll_{i}(\sqrt{\mathcal{Y}\ell}/c)^{i}$ for any $i\geq 0$. The next move is to resort to Lemma 2.3; this would yield the following estimate
$\ll\sup_{C^{\star}\leq C\ll\sqrt{\mathcal{Y}\ell}/K^{2-\varepsilon}}\sum_{c\sim C}\frac{\sqrt{\mathcal{Y}\ell}}{c^{3/2}}\cdot\mathcal{Y}^{1/8+\varepsilon}\ell^{3/8}\ll\frac{\mathcal{Y}^{5/8+\varepsilon}\ell^{7/8}}{\sqrt{C^{\star}}},$ (4.69)
where it is necessary that $\ell\gg K^{4}/X$ and $C^{\star}<\sqrt{\mathcal{Y}\ell}/K^{2-\varepsilon}$ to ensure that the sum over $c$ is not vacuous. Combining with (4.63), it follows that
$\mathcal{T}(K,t)\ll\mathcal{Y}^{9/16+\varepsilon}\ell^{7/16},$
upon choosing $C^{\star}=\mathcal{Y}^{1/8}\ell^{7/8}$. One observes that $C^{\star}<\sqrt{\mathcal{Y}\ell}/K^{2-\varepsilon}$, provided that $\ell<\mathcal{Y}^{1-\varepsilon}/K^{16/3}$.
4.1.2. Bounding $\widehat{\mathcal{T}}$
Next, in this part let us take a look at the term $\widehat{\mathcal{T}}$ in (4.59). This can be handled by a similar argument. Actually, by repeating the procedure as in §4.1, one sees that $\widehat{\mathcal{T}}$ is boiled down to evaluating
$\begin{split}\sup_{\sqrt{\mathcal{Y}\ell}/K^{1-\varepsilon}\ll C\ll\sqrt{\mathcal{Y}\ell}/K^{1-100\varepsilon}}\frac{C}{\sqrt{\mathcal{Y}}}&\sum_{c\sim C}\sum_{\begin{subarray}{c}\pm,m\\ m\geq 1\end{subarray}}\frac{\overline{A_{F}(m,1)}}{m}\,e\left(\mp\frac{m\overline{\ell}}{c}\right)\,\Omega_{\pm}\left({\frac{m\mathcal{Y}}{c^{3}};{U}_{c,\ell}}\right),\end{split}$ (4.70)
with ${U}_{c,\ell}(x)=V(x)e{\left(-{t\log x\mathcal{Y}}/{2\pi}\right)}$ for any $x>0$. By Lemma 2.2, the display above turns out to be controlled by
$\frac{\mathcal{Y}^{1/6}}{C}\sum_{c\sim C}\sum_{\begin{subarray}{c}\pm,m\geq 1\end{subarray}}\frac{\overline{A_{F}(m,1)}}{{m}^{1/3}}e\left(\mp\frac{m\overline{\ell}}{c}\right)\,\widehat{U_{\pm,\ell}}\left({\frac{m\mathcal{Y}}{c^{3}}}\right),$ (4.71)
where $\widehat{U_{\pm,\ell}}(y)=\int_{\mathbb{R}^{+}}V(x)e(-t\log x/2\pi\pm 3(yx)^{1/3})\mathrm{d}x$ for any $y\in\mathbb{R}^{+}$. Recall that $t\gg\mathcal{Y}^{100\varepsilon}$. The ‘+’ case only survives. In that case, the integral is truncated at $x\asymp|t|^{3}$, and is roughly given by $x^{-1/6}e(\varrho t)V(t^{3}/(8\pi^{2}x))$ for some non-zero constant $\varrho\in\mathbb{R}$. This reveals that the display above is dominated by the upper-bound for
$\begin{split}&\frac{1}{\sqrt{C}}\sum_{c\sim C}\,\sum_{m\asymp C^{3}|t|^{3}/\mathcal{Y}}\frac{\overline{A_{F}(m,1)}}{\sqrt{m}}\,e\left(\frac{m\overline{\ell}}{c}\right)\,V^{\ddagger}\left(\frac{8\pi^{2}m\mathcal{Y}}{C^{3}|t|^{3}}\right),\end{split}$ (4.72)
where $V^{\ddagger}(x)=e{\left(\varrho/(2\pi^{2/3})tx^{1/3}\right)}V(x)$ for any $x\in\mathbb{R}^{+}$, satisfying that $\frac{\partial^{i}}{\partial x^{i}}{V^{\ddagger}}(x)\ll_{j}|t|^{j}$ for any $j\geq 0$. Notice that, here, if $\ell\gg K^{1+\varepsilon}/|t|^{3}$, one has $C^{3}|t|^{3}/\mathcal{Y}\gg 1$. Now, we attempt to apply Lemma 2.3, which shows the estimate that
$\widehat{\mathcal{T}}(K,t)\ll\frac{|t|^{7/4}C^{5/4}}{\mathcal{Y}^{1/4-\varepsilon}}\ll\frac{|t|^{7/4}\mathcal{Y}^{3/8+\varepsilon}\ell^{5/8}}{K^{5/4}}.$ (4.73)
In summary, upon collecting the bounds in (4.60), (4.61), (4.69) and (4.73), we obtain
Proposition 4.2.
For any $\ell\in\mathbb{R}^{+}$ such that
$1+\frac{K^{4+\varepsilon}}{\mathcal{Y}}+\frac{K^{1+\varepsilon}}{|t|^{3}}\leq\ell<\frac{\mathcal{Y}^{1-\varepsilon}}{K^{16/3}},$ (4.74)
with $K$ sufficiently large and $K^{\varepsilon}\ll|t|\ll K^{2-\varepsilon}$, we have

### v4

@@ 3.1.2. Oscillatory case
$\begin{split}&L\geq\frac{M^{1+\varepsilon}}{\sqrt{\mathcal{X}}},\quad L\leq\frac{\mathcal{X}^{1/12-\varepsilon}}{q^{5/12}}\\ &L\leq\frac{1}{\sqrt{q}}\left(\frac{\mathcal{X}}{M}\right)^{1/4-\varepsilon},\quad L\leq\left(\frac{\mathcal{X}}{q^{2}}\right)^{1/32-\varepsilon};\end{split}$ (3.44)
whilst, the second line requires that
$\begin{split}&L\leq\frac{\mathcal{X}^{1/28-\varepsilon}M^{1/14}}{q^{3/7}},\quad L\leq\left(\frac{\mathcal{X}}{q^{6}}\right)^{1/16-\varepsilon}\quad L\leq\frac{\mathcal{X}^{1/20-\varepsilon}M^{1/10}}{q^{2/5}}.\end{split}$ (3.45)
Recall that $(M^{3/2}q^{2})^{1-\delta}\leq\mathcal{X}\leq(M^{3/2}q^{2})^{1+\varepsilon}$ for any small constants $\delta,\varepsilon>0$. Collecting all the inequalities (3.1), (3.2), (3.37), (3.44) and (3.45), we infer that the effective range for $L$ should be
$\frac{q^{1+\varepsilon}}{{M^{1/4}}}+\frac{M^{(1+3\delta)/4+\varepsilon}}{q^{1-\delta}}+1\leq L\leq\min{\left\{M^{3/64-\varepsilon},\frac{M^{3/32-\varepsilon}}{{q}^{1/4}},\,\frac{{M}^{1/8}}{q^{5/14+\varepsilon}},\,\frac{M^{7/40-\varepsilon}}{q^{3/10}}\right\}}.$ (3.46)
This yields the range $M^{13/64+\varepsilon}\leq q\leq M^{11/40-\varepsilon}$, and thus gives the requirement as shown in (1.4). The assertion of Corollary 1.3 now follows by combining with (3.2).
4. Proofs of Theorems 1.4 & 1.5
In this section, we shall be dedicated to the proofs of Theorem 1.4 and Theorem 1.5.
@@
We shall consider the average of the first moments of $\rm\text{GL}(3)\times GL(2)$ and $\text{GL}(2)$ $L$-functions in a family with a ‘harmonic’ average of $k\sim K$. To be precise, let $W$ be a smooth function compactly supported on $[1/2,5/2]$ with bounded derivatives. Let $f$ be a $\text{GL}(2)$-newform of weight $k\in 2\mathbb{Z}$ and full level. Let $K$ be a sufficiently large parameter. Let $\ell\geq 1$ be an integer such that $K^{1-\delta}\ll\ell\ll K^{1+\varepsilon}$ for any $\delta,\varepsilon>0$. We attempt to obtain an upper-bound for
$\sum_{k\equiv 0\bmod 2}W{\left(\frac{k-1}{K}\right)}\sum_{f\in\mathcal{B}^{\ast}_{k}}\omega^{-1}_{f}\lambda_{f}(\ell)L(1/2,F\otimes f).$ (4.1)
To begin with, we will give a precise expression for the central value of $L(s,F\otimes f)$. This can be derived in a standard way from [23, Theorem 5.3]. Indeed, analogous to [18, Eqn. (4.11)], one finds that
$L(1/2,F\otimes f)=(1+\varepsilon(F\otimes f))\sum_{m,n\geq 1}\frac{\mu(m)A_{F}(n,1)\lambda_{f}(mn)}{\sqrt{m^{3}n}}V^{\star}\left(\frac{m^{3}n}{k^{3}}\right),$ (4.2)
where
$V^{\star}(y)=\frac{1}{2\pi i}\int_{(3)}\left(\cos\left(\frac{\pi s}{4A}\right)\right)^{-24A}\frac{L_{\infty}(s+1/2,F\otimes f)}{L_{\infty}(1/2,F\otimes f)}L(1+2s,F)y^{-s}\frac{\mathrm{d}s}{s},$
with $A>0$ sufficiently large. Upon shifting the line to $\text{Re}(s)=-1/2$, we get
$V^{\star}(y)=L(1,F)+O(k^{\varepsilon}y^{1/2+\varepsilon})$ (4.3)
for $y\leq 1$. Whilst, for $y>1$, one might shift the line to $\text{Re}(s)=A$, obtaining
$y^{j}{V^{\star}}^{(j)}(y)\ll\left(\frac{k^{3}}{y}\right)^{A}$
for any $j\geq 0$, upon exploiting Stirling’s estimates to (2.5). As described in §2.3, the root number $\varepsilon(F\otimes f)$ is dependent on the weight $k$; thus, by (4.2), we need to turn to seeking an effective control over the following two sums
$\sum_{k\equiv 0\bmod 2}W{\left(\frac{k-1}{K}\right)}\sum_{f\in\mathcal{B}^{\ast}_{k}}\omega^{-1}_{f}\lambda_{f}(\ell)\,\sum_{n\geq 1}\frac{A_{F}(n,1)\lambda_{f}(n)}{\sqrt{n}}V^{\star}{\left(\frac{n}{\mathcal{Y}}\right)},$ (4.4)
@@
$\sum_{k\equiv 0\bmod 2}i^{k}\,W{\left(\frac{k-1}{K}\right)}\sum_{f\in\mathcal{B}^{\ast}_{k}}\omega^{-1}_{f}\lambda_{f}(\ell)\,\sum_{n\geq 1}\frac{A_{F}(n,1)\lambda_{f}(n)}{\sqrt{n}}V^{\star}{\left(\frac{n}{\mathcal{Y}}\right)},$ (4.5)
where $k^{3-\delta}\leq\mathcal{Y}\leq k^{3-\varepsilon}$ for arbitrarily small constants $\delta,\varepsilon>0$. By equipping with Petersoon formula, Lemma 2.4, we split the expressions in (4.5) and (4.5) into the diagonal terms and the off-diagonal terms. Via (4.3), both diagonal terms add up to
$\sum_{k\equiv 0\bmod 2}(1+i^{k})\,W{\left(\frac{k-1}{K}\right)}\frac{A_{F}(\ell,1)\lambda_{f}(\ell)}{\sqrt{\ell}}V^{\star}{\left(\frac{\ell}{\mathcal{Y}}\right)}=$ $\frac{L(1,F)A_{F}(\ell,1)\lambda_{f}(\ell)K}{4\sqrt{\ell}}\widehat{W}(0)$
$+O(K^{-1/2+\varepsilon}),$ (4.6)
with $\widehat{W}$ being the Fourier transform of $W$, which is given by $\widehat{W}(x)=\int_{\mathbb{R}}W(y)e(xy)\mathrm{d}y$ for any $x\in\mathbb{R}^{+}$.
We shall merely handle the off-diagonal term for the second sum (4.5) in the following. The same analysis works for the first sum (4.4). The main difference is that, for the first sum, the Lemma 2.8 will be used, which, however, produces a possibly weaker bound since of more error terms thereof; for the second one, we will apply Lemma 2.9 instead. It is verifiable that the off-diagonal term for (4.5) is presented as the following form
$2\pi\sum_{c\geq 1}\sum_{k\equiv 0\bmod 2}W{\left(\frac{k-1}{K}\right)}\sum_{n\geq 1}\frac{A_{F}(n,1)}{\sqrt{n}}\frac{S(n,\ell;c)}{c}J_{k-1}{\left(\frac{4\pi\sqrt{n\ell}}{c}\right)}V^{\star}{\left(\frac{n}{\mathcal{Y}}\right)}.$
@@
$\mathcal{T}(K,\ell)=\sum_{c\ll\sqrt{\mathcal{Y}\ell}/K^{2-\varepsilon}}\,\sum_{n\geq 1}\frac{A_{F}(n,1)}{\sqrt{n}}\frac{S(n,\ell;c)}{c}\,e{\left(\frac{2\sqrt{n\ell}}{c}\right)}V^{\star}{\left(\frac{n}{\mathcal{Y}}\right)},$ (4.7)
@@
$\widehat{\mathcal{T}}(K,\ell)=\sum_{\sqrt{\mathcal{Y}\ell}/K^{1-\varepsilon}\ll c\ll\sqrt{\mathcal{Y}\ell}/K^{1-100\varepsilon}}\,\sum_{n\geq 1}\,\frac{A_{F}(n,1)}{\sqrt{n}}\frac{S(n,\ell;c)}{c}W{\left(\frac{4\pi\sqrt{n\ell}}{cK}\right)}V^{\star}{\left(\frac{n}{\mathcal{Y}}\right)}$ (4.8)
plus two error terms $\text{Err}^{\flat}(K,\ell))$ and $\text{Err}^{\natural}(K,\ell)$, say, which can be estimated as
$\begin{split}\text{Err}^{\flat}(K,\ell))&\ll\frac{1}{K^{2-\varepsilon}}\sum_{n\sim\mathcal{Y}}\frac{|A_{F}(n,1)|}{\sqrt{n}}\sum_{c\sim\sqrt{\mathcal{Y}\ell}/K}\frac{|S(n,\ell;c)|}{c}\ll\frac{\mathcal{Y}^{3/4}\ell^{1/4}}{K^{5/2-\varepsilon}},\end{split}$ (4.9)
@@
$\begin{split}\text{Err}^{\natural}(K,\ell))&\ll\frac{\sqrt{\ell}}{K^{4-\varepsilon}}\sum_{n\sim\mathcal{Y}}{|A_{F}(n,1)|}\sum_{c\ll\sqrt{\ell}\mathcal{Y}^{1/2+\varepsilon}}\frac{|S(n,\ell;c)|}{c^{2}}\ll\frac{\mathcal{Y}\sqrt{\ell}}{K^{4-\varepsilon}},\end{split}$ (4.10)
respectively. Here, $\text{Err}^{\flat}(K,\ell)$ comes from the first error term in (2.19); while, $\text{Err}^{\natural}(K,\ell)$ is for the second one. We have applied bound (2.14) in estimating $\text{Err}^{\natural}(K,\ell)$, when $k$ is sufficiently large.
[…]
${\mathcal{T}}(K,\ell)\ll\frac{1}{\mathcal{Y}^{1/2-\varepsilon}}\left(1+(\mathcal{Y}\ell)^{1/2+\varepsilon}\right)\ll\ell^{1/2+\varepsilon}.$ (4.11)
Next, let us take a look at the term $\widehat{\mathcal{T}}$ in (4.8). By invoking the Voronoĭ summation formula, Lemma 2.1, this sum is thus essentially transformed into
$\begin{split}\sup_{\sqrt{\mathcal{Y}\ell}/K^{1-\varepsilon}\ll C\ll\sqrt{\mathcal{Y}\ell}/K^{1-100\varepsilon}}\frac{C}{\sqrt{\mathcal{Y}}}&\sum_{c\sim C}\sum_{\begin{subarray}{c}\pm,m^{\ast},r\\ m^{\ast}\geq 1\\ r|c\end{subarray}}\frac{\overline{A_{F}(m^{\ast},r)}}{m^{\ast}r^{2}}\,e\left(\mp\frac{m^{\ast}r^{2}\overline{\ell}}{c}\right)\,\Omega_{\pm}\left({\frac{m^{\ast}r^{2}\mathcal{Y}}{c^{3}};{V}_{c,\ell}}\right),\end{split}$
with $V_{c,\ell}(x)=V^{\star}(x)W{({4\pi\sqrt{x\mathcal{Y}\ell}}/{(cK)})}$ for any $x\in\mathbb{R}^{+}$. By Lemma 2.2, the display above turns out to be controlled by
$\begin{split}\sup_{\sqrt{\mathcal{Y}\ell}/K^{1-\varepsilon}\ll C\ll\sqrt{\mathcal{Y}\ell}/K^{1-100\varepsilon}}\frac{\mathcal{Y}^{1/6}}{C}&\sum_{c\sim C}\sum_{\begin{subarray}{c}\pm,m^{\ast},r\\ m^{\ast}\geq 1\\ r|c\end{subarray}}\frac{\overline{A_{F}(m^{\ast},r)}}{{m^{\ast}}^{1/3}r^{2/3}}\,e\left(\mp\frac{m^{\ast}r^{2}\overline{\ell}}{c}\right)\,\widehat{V_{\pm,\ell}}\left({\frac{m^{\ast}r^{2}\mathcal{Y}}{c^{3}}}\right),\end{split}$ (4.12)
@@
$\widehat{V_{\pm,\ell}}(x)=\int_{\mathbb{R}^{+}}V^{\star}(y)W{\left(\frac{4\pi\sqrt{y\mathcal{Y}\ell}}{cK}\right)}\,e{\left(\pm 3(xy)^{1/3}\right)}\mathrm{d}y.$
A direct computation with the help of the Kim–Sarnak bound shows an upper-bound of $O(\mathcal{Y}^{1/6+\varepsilon})$ for (4.12); this, however, is sufficient for our purpose. In summary, upon collecting the estimates in (4.1), (4.9), (4.10) and (4.11), we obtain